# Efficient IoT Intrusion Detection with an Improved Attention-Based CNN-BiLSTM Architecture

### v3

@@ Abstract
The ever-increasing security vulnerabilities in the Internet-of-Things (IoT) systems require improved threat detection approaches. This paper presents a compact and efficient approach to detect botnet attacks by employing an integrated approach that consists of traffic pattern analysis, temporal support learning, and focused feature extraction. The proposed attention-based model benefits from a hybrid CNN-BiLSTM architecture and achieves 99% classification accuracy in detecting botnet attacks utilizing the N-BaIoT dataset, while maintaining high precision and recall across various scenarios. The proposed model’s performance is further validated by key parameters, such as Mathews Correlation Coefficient and Cohen’s kappa Correlation Coefficient. The close-to-ideal results for these parameters demonstrate the proposed model’s ability to detect botnet attacks accurately and efficiently in practical settings and on unseen data. The proposed model proved to be a powerful defence mechanism for IoT networks to face emerging security challenges.
Index Terms:
Botnets, intrusion detection, BiLSTM, IoT, attention mechanism.
@@ I Introduction
The application of Artificial Intelligence (AI) in Internet-of-Things (IoT) has resulted in innovative and sophisticated networked systems. Such systems, comprising of a number of devices ranging from simple sensors to advanced machinery, influence a number of applications, e.g., smart cities, homes, transportation, manufacturing, and agriculture [1]. IoT devices have become deeply embedded in everyday life. However,this extensive adoption brings significant security risks. Due to their diverse nature and inherent vulnerabilities, IoT devices are particularly prone to malicious activities or cyber attacks such as malware attacks. As the use of IoT devices continues to grow, concerns over security and privacy are rising, given the increasing exposure of sensitive data online. To reduce these cyber attacks, continuous network monitoring is crucial. Analysing IoT network traffic can enhance intrusion detection mechanisms within the IoT ecosystem, ultimately strengthening the cybersecurity of networks [2]. Intrusion detection systems serve as essential supplementary security measures, operating in conjunction with other defences to reduce potential threats. Deep Learning, a subset of machine learning, is capable of extracting relevant information from massive and high-speed network traffic generated by various interconnected IoT devices [3].
This paper introduces a hybrid deep learning security framework that integrates an attention mechanism with 1D convolutional neural networks (CNN) and Bidirectional Long Short-Term Memory (BiLSTM) architectures. The framework demonstrates robust intrusion detection capabilities using normalized data and measuring performance using the N-BaIoT dataset across various performance and reliability metrics, taking into account the specific challenges of the datasets, including device behaviours and class imbalance [4]. This study presents the following key contributions:
  1. 1.
    A hybrid attention-based deep learning model that integrates 1D-CNN + BiLSTM architectures to efficiently and accurately detect botnet attacks.
  2. 2.
@@ II Literature Review
The increase in the number of resource-constrained IoT devices has made their security more challenging, as novel security risks emerge. Recently, researchers have been developing a lot of cryptography techniques [5, 6] and artificial intelligence-based IDS to protect IoT systems. A number of machine learning methods can be found that utilize N-BaIoT data to address these security concerns, for example, in [7], the authors presented a collaborative learning model for an improved IDS intended to secure IoT industry 4.0. Their methodology, which merge cooperative learning and Deep Neural Networks (DNN) approaches, scored well, obtaining an 0.92% as F1-Score and a 95.06% of accuracy. Another study [8], utilising several neural network structures demonstrated that CNN, RNN and LSTM outperformed in botnet identification with an accuracy of 91%, 41% and 62% respectively. Moreover, the authors in [9] developed a CNN-based deep learning model for botnet detection using the N-BaIoT dataset, improving scalability in diverse environments. The model showed superior accuracy, precision, and recall, but struggled with UDP-based attacks. The research emphasizes the effectiveness of CNNs in IoT security. Similarly, a novel CNN-LSTM model has been introduced in [10], which detects BASHLITE and Mirai security attacks, achieving an accuracy of 90.88% and precision and recall rates of 93.04% and 91.91% respectively. This work marks a significant improvement in attack detection capabilities. These studies demonstrate the advancement and efficacy of several algorithms in addressing the challenges connected with IoT botnet detection.
@@ III Preliminaries
The proposed model integrates 1D-CNN, BiLSTM and an attention mechanism to create a robust architecture for sequential data analysis. This hybrid approach capitalizes on CNN’s ability to detect local patterns and spatial relationships, BiLSTM’s capacity to capture long-range dependencies, and the attention mechanisms focus on crucial input elements.
III-A 1D-CNNs
1D-CNNs have shown particularly effective for classification methods, adopting a multi-layered design that consists of five different layers for processing.The Convolutional layers serve a substantial part in determining feature extraction and parameter optimization during back propagation. These structures make use of dedicated filters, incorporating pooling techniques to prevent over-fitting. The data is then turned into class representations using flattening and connected layers. The modular design provides structural flexibility to meet a variety of categorization requirements. The mathematical equation[11] of the 1D – CNN are as follows:
@@ III-B Bidirectional Long Short-Term Memory (Bi-LSTM)
Bi-LSTM analyses sequences bidirectionally to detect connections and temporal patterns. Its nature of two-way analysis improves its ability to recognize complicated sequential patterns.
Figure 1: The BiLSTM Block [12].
@@ III-C Attention Mechanism
$\mathbf{K}=\tanh(\mathbf{VW}^{a})$ (5)
$\mathbf{d}=\text{softmax}(\mathbf{qK}^{T}),\quad\mathbf{a}=\mathbf{dV}$ (6)
@@ V-A Dataset
The N-BaIoT dataset [14]incorporates 7,062,606 network traffic data patterns from several smart devices to identify attacks like botnet. It comprises of 115 features determined across many time frames, involving both valid and malicious activities on the network. It includes traffic data from nine business IoT devices infected with malware either Mirai or BASHLITE,conveying 10 various attack types. The dataset has 10 classes of attacks with 1 class as benign.
@@ V-B1 Performance Parameters
To analyze the performance of the architecture, we used standard metrics resulting from the confusion matrix. These indicators comprise of accuracy, recall, precision, and F1-score all these provides a dept insight into the model’s classification capabilities. Four key elements are involved in confusion matrix: False Positives (FP)—Normal instances mistakenly classified as attacks, False Negatives (FN)—Attack instances mistakenly classified as normal, True Positives (TP)—Correctly identified attack instances, True Negatives (TN)—Correctly identified normal instances. Using these elements, we calculate the following indicators:
1. Accuracy: The percentage of accurate predictions (which involves true positives and true negatives) across all instances examined.
@@ VIII Conclusion
This paper proposed and evaluated a deep learning model for botnet attack detection, which incorporated an attention process into a hybrid 1D-CNN and Bi-LSTM architecture. The proposed model demonstrated 99% accuracy and close to ideal results for reliability parameters including Matthews Correlation coefficient and Cohen’s kappa coefficient. Moreover, the optimized architecture ensured minimal computation costs, making it perfect for real-time monitoring and protection of IoT devices.

### v4

@@ Abstract
The ever-increasing security vulnerabilities in the Internet-of-Things (IoT) systems require improved threat detection approaches. This paper presents a compact and efficient approach to detect botnet attacks by employing an integrated approach that consists of traffic pattern analysis, temporal support learning, and focused feature extraction. The proposed attention-based model benefits from a hybrid CNN-BiLSTM architecture and achieves 99% classification accuracy in detecting botnet attacks utilizing the N-BaIoT dataset, while maintaining high precision and recall across various scenarios. The proposed model’s performance is further validated by key parameters, such as Mathews Correlation Coefficient and Cohen’s kappa Correlation Coefficient. The close-to-ideal results for these parameters demonstrate the proposed model’s ability to detect botnet attacks accurately and efficiently in practical settings and on unseen data. The proposed model proved to be a powerful defense mechanism for IoT networks to face emerging security challenges.
@@ I Introduction
The application of Artificial Intelligence (AI) in Internet-of-Things (IoT) has resulted in innovative and sophisticated networked systems. Such systems, comprising of a number of devices ranging from simple sensors to advanced machinery, influence a number of applications, e.g., smart cities, homes, transportation, manufacturing, and agriculture [8]. IoT devices have become deeply embedded in everyday life. However, this extensive adoption brings significant security risks. Due to their diverse nature and inherent vulnerabilities, IoT devices are particularly prone to malicious activities or cyber attacks such as malware attacks. As the use of IoT devices continues to grow, concerns over security and privacy are rising, given the increasing exposure of sensitive data online. To reduce cyber attacks targeting IoT networks, continuous monitoring is crucial. Analyzing IoT network traffic can enhance intrusion detection mechanisms within the IoT ecosystem,ultimately strengthening cybersecurity [4]. Intrusion detection systems serve as essential supplementary security measures, operating in conjunction with other defenses to reduce potential threats. Deep Learning a subset of machine learning is capable of extracting relevant information from massive and high-speed network traffic generated by various interconnected IoT devices [11].The presented hybrid deep learning security framework integrates the attention mechanism with 1D-CNN and BiLSTM. The framework demonstrates robust intrusion detection capabilities using normalized data and measuring performance using the N-BaIoT dataset across common metrics. Common metrics are used to evaluate the performance of each model, taking into account the specific challenges of the data sets, including various device behaviors and class imbalance [5] This study presents a few significant advancements which are as follows:
  • The proposed attention-based 1D-CNN + BiLSTM model efficiently and accurately detects botnet attacks while maintaining high-precision and recall in various scenarios.
@@ II Literature Review
The growing number of limited resources IoT devices is becoming more harder to protect as novel security risks emerge. Researchers have been practicing with a number of machine learning methods utilizing N-BaIoT data to figureout these security concerns, for instance, authors in [7] demonstrated an IDS for IoT industry 4.0. They merged cooperative learning and Deep Neural Networks (DNN) approaches and the proposed scheme scored well, obtaining an 0.92% as F1-Score and a 95.06% of accuracy. Another study [12], on several neural network structures, discovered that CNN, RNN and LSTM outperformed in botnet identification with an accuracy of 91%, 41% and 62%, respectively. Furthermore, the researchers in [9] developed a CNN-based deep learning model for botnet detection using the N-BaIoT dataset, improving scalability in diverse environments. The model showed superior accuracy, precision, and recall, but struggled with UDP-based attacks. The research emphasizes the effectiveness of CNNs in IoT security. In addition, a novel CNN-LSTM model [1] was introduced to identify BASHLITE and Mirai security attacks, achieving an accuracy of 90.88% and precision and recall rates of 93.04% and 91.91%, respectively. This work marks a significant improvement in attack detection capabilities. These studies demonstrate the advancement and efficacy of several algorithms in addressing the challenges connected with IoT botnet detection.
@@ III Preliminaries
The proposed model integrates 1D-CNN’s, BiLSTM and an attention mechanism to create a robust architecture for sequential data analysis. This hybrid approach capitalizes on CNN’s ability to detect local patterns and spatial relationships, BiLSTM’s capacity to capture long-range dependencies, and the attention mechanisms focus on crucial input elements. By incorporating two complimentary techniques, the model aims to boost interpretability and performance across a range of sequential data requirements, providing additional and sophisticated analysis than common DL strategies.
III-A 1D-CNN’s
One-Dimensional Convolutional Neural Network has shown particularly effective for classification methods, adopting a multi-layered design that consists of five different layers for processing.The Convolutional layers serve a substantial part in determining feature extraction and parameter optimization during back propagation. These structures make use of dedicated filters, incorporating pooling techniques to prevent over-fitting. The data is then turned into class representations using flattening and connected layers. The modular design provides structural flexibility to meet a variety of categorization requirements. The mathematical equation[6] of the 1D – CNN are as follows:
@@ III-B Bidirectional Long Short-Term Memory (Bi-LSTM)
Bi-LSTM analyzes sequences bidirectionally to detect connections and temporal patterns. Its nature of two-way analysis improves its ability to recognize complicated sequential patterns.
Figure 1: BiLSTM [14].
@@ III-C Attention Mechanism
$\mathbf{K}$ $=\tanh(\mathbf{V}\mathbf{W}^{a})$ (5)
$\mathbf{d}$ $=\operatorname{softmax}(\mathbf{q}\mathbf{K}^{T}),\quad\mathbf{a}=\mathbf{d}\mathbf{V}$ (6)
@@ V-A Dataset
The N-BaIoT dataset[10] incorporates 7,062,606 network traffic data patterns from several smart devices to identify attacks like botnet. It comprises of 115 features determined across many time frames, involving both valid and malicious activities on the network. It includes traffic data from nine business IoT devices infected with malware either Mirai or BASHLITE,conveying 10 various attack types. This rich dataset provides a robust foundation for developing and evaluating intrusion detection systems in IoT networks. The dataset has 10 classes of attacks with 1 class as benign.
@@ V-B1 Performance Parameters
To analyze the performance of the architecture, we use standard metrics resulting from the confusion matrix. These indicators comprise of accuracy, recall, precision, and F1-score all these provides a dept insight into the model’s classification capabilities. Four key elements are involved in confusion matrix:
  1. 1.
    False Positives (FP): Normal instances mistakenly classified as attacks.
  2. 2.
    False Negatives (FN): Attack instances mistakenly classified as normal.
  3. 3.
    True Positives (TP): Correctly identified attack instances.
  4. 4.
    True Negatives (TN): Correctly identified normal instances.
Using these elements, we calculate the following indicators: 1. Accuracy: The percentage of accurate predictions (which involves true positives and true negatives) across all instances examined.
@@ VIII Conclusion
This paper proposed and evaluated a deep learning model for botnet attack detection, which incorporates an attention process into a hybrid 1D- convolutional and Bi-LSTM architecture. The proposed model demonstrated 99% accuracy and close to ideal results for reliability parameters including Matthews Correlation coefficient and Cohen’s kappa coefficient. Moreover, the optimized architecture ensured minimal computation costs, making it perfect for real-time monitoring and protection of IoT devices.